What mass of calcium chloride will be produced from 4.09 g of HCl reacting with Ca(OH)2?
1
Expert's answer
2020-04-27T01:53:05-0400

Ca(OH)2+2HCl=CaCl2+2H2O


n(HCl)=m(HCl)/M(HCl)=4.09/36.5=0.11mol

n(HCl):n(CaCl2)=2:1

n(CaCl2)=0.11/2=0.056mol


m(CaCl2)=n(CaCl2)*M(CaCl2)=0.056*111=6.216g


Answer: m(CaCl2)=6.216g